Given the following regression equation for predicting the expected burglary rate for a given level of poverty, Y = 145.23 + 12.56X, determine the projected burglary rates for:

  1. 25% poverty
  1. 50% poverty

  1. 75% poverty
Answer #1

Solution:

Given

Y=145.23+12.56X

Estimate the projected burglary rates for given level of poverty.

Part A) X=25%

Thus

Y = 145.23+12.56X

Y = 145.23+12.56*25%

Y = 145.23+3.14

Y= 148.37

Part B) X=50%

Y = 145.23+12.56*50%

Y = 145.23 + 6.28

Y = 151.51

Part C) X=75%

Y = 145.23+12.56*75%

Y = 145.23+9.42

Y = 154.65